### v1

# On Lev’s periodicity conjecture

Abstract.
We classify the sum-free subsets of ${\mathds{F}}_{3}^{n}$ whose density exceeds $\frac{1}{6}$. This yields a resolution of Vsevolod Lev’s periodicity conjecture, which asserts that if a sum-free subset ${A\subseteq{\mathds{F}}_{3}^{n}}$ is maximal with respect to inclusion and aperiodic (in the sense that there is no non-zero vector $v$ satisfying $A+v=A$), then $|A|\leq\frac{1}{2}(3^{n-1}+1)$ —a bound known to be optimal if $n\neq 2$, while for $n=2$ there are no such sets.
Key words and phrases:
sum-free sets, ternary vector spaces, Lev’s periodicity conjecture
2010 Mathematics Subject Classification:
11B13, 11B30, 11P70
1. Introduction
A subset $A$ of an abelian group $G$ is said to be sum-free if the equation $x+y=z$ has no solutions with $x,y,z\in A$. Thus Fermat’s last theorem is equivalent to the statement that for every $k\geq 3$ the set of positive $k^{\mathrm{th}}$ powers is sum-free; this motivated Schur [Schur] to prove his famous theorem that the set of positive integers cannot be represented as a union of finitely many sum-free sets. Since the 1960’s the study of sum-free sets has intensified due to the problems and results of Erdős (see e.g. [E65]). We refer to [Green, EGM] for some recent milestones in the area and to [TV17] for a survey.
When the ambient group is a binary vector space $G={\mathds{F}}_{2}^{n}$, sum-free sets have connections to projective geometry and coding theory. It is easily seen that in this case the sum-free sets of largest cardinality are hyperplanes not passing through the origin. Davydov and Tombak [DT] obtained the much deeper result that sum-free sets which cannot be covered by a hyperplane have size at most $5\cdot 2^{n-4}$. They also have a stronger theorem on maximal sum-free sets in ${\mathds{F}}_{2}^{n}$, i.e., sum-free sets that are maximal with respect to inclusion. Notably they showed that every such set $A$ possessing more than $12^{n-2}+$ 1 elements is periodic, which means that there is a non-zero vector $v$ such that $A+v=A$.
Vsevolod Lev [VL05, VL23] initiated the study of similar questions in other finite vector spaces. Following the title of [VL05] we call vector spaces over the three-element field ${\mathds{F}}_{3}$ ternary. By an early observation of Wallis, Street, and Wallis [WSW]*Corollary 7.11 sum-free subsets of ${\mathds{F}}_{3}^{n}$ have at most the size $3^{n-1}$, the extremal examples again being hyperplanes not passing through the origin. The main result of [VL05] asserts that if a sum-free subset of ${\mathds{F}}_{3}^{n}$ has more than $5\cdot 3^{n-3}$ elements, then it can be covered by a hyperplane; this bound is indeed optimal whenever ${n\geq 3}$. Furthermore, Lev conjectured in [VL05] that the largest possible cardinality of an aperiodic sum-free subset of ${\mathds{F}}_{3}^{n}$ is $\frac{1}{2}(3^{n-1}+1)$ (when $n\neq 2$) and provided a construction exemplifying the lower bound. Despite being included as Problem 7.10 into the well-known problem collection of Croot and Lev [CL07], the conjecture has remained open for almost two decades.
Before we formulate our main result, which classifies all sum-free subsets of ${\mathds{F}}_{3}^{n}$ whose cardinality is at least $\frac{1}{2}(3^{n-1}+1)$, we would like to fix some notation concerning affine subspaces. Given an affine subspace $U$ of a vector space $V$ we write $[U]=U-U$ for the translate of $U$ passing through the origin, and $\dim(U)=\dim([U])$ for its dimension. When we have two affine subspaces $U\subseteq H\subseteq V$, then $H/U=H/[U]$ means the set of translates of $U$ comprising $H$ and $\dim(H/U)$ is shorthand for $\dim(H)-\dim(U)$, i.e., the dimension of the affine subspace $H/[U]$ of the quotient space $V/[U]$. As usual, $H$ is called a hyperplane if $\dim(V/H)=1$. The affine space generated by an arbitrary set $X\subseteq V$ is designated by $\mathrm{aff}(X)$. Finally, if $U$ is an affine subspace of $V$, then the vector subspace of $V$ generated by $U$ is called the cone over $U$ and denoted by $C(U)$. For instance, if $V$ is ternary and $0\not\in U$, then $C(U)$ is the disjoint union of $[U]$, $U$, and $-U$.
Definition 1.1.
Given two affine subspaces $U\subseteq H$ of a ternary vector space we call a set $W\subseteq H$ an $(H,U)$-half if $W+[U]=W$ and $H$ is the disjoint union of $U$, $W$, and $(-U)+(-W)$.
Notice that the condition $W+[U]=W$ just means that $W$ is a union of translates of $U$. Moreover, $H/U$ consists of $U$ itself and $\frac{1}{2}(|H/U|-1)$ pairs $\{U+x,U-x\}$; the second demand on $W$ is equivalent to $W$ containing exactly one translate from each such pair. For this reason, all $(H,U)$-halves $W$ have the same cardinality $|W|=\frac{1}{2}(|H|-|U|)$. We are now ready for the central definition of this article, which proceeds by recursion on the dimension of the ambient ternary space $V$.
Definition 1.2.
Given a finite ternary space $V$ we call a subset $A\subseteq V$ primitive if
  1. ()
    either $A$ is a hyperplane not containing the origin,
  2. ()
    or there are a hyperplane $H\subseteq V$ not passing through the origin, a proper affine subspace ${U\subsetneqq H}$, an $(H,U)$-half $W$, and a primitive set $X\subseteq C(U)$ such that
    1. ()
      $A=W\cup X$,
    2. ()
      $X\cap[U]=\varnothing$,
    3. ()
      $\dim(H/U)\geq 2$ or $X\neq-U$,
    4. ()
      and $\mathrm{aff}(X\cap(-U))=-U$.
As it turns out, primitive sets are the same as dense maximal sum-free sets. Their definition might nevertheless appear complicated on first reading, and thus we would like to discuss a few low-dimensional examples before moving on. In ${\mathds{F}}_{3}$ itself the only primitive sets are $\{1\}$ and $\{2\}$. Assume next that there existed a primitive subset of ${\mathds{F}}_{3}^{2}$ falling under clause ()1.2. Without loss of generality $H$ is the line $\{(1,y)\colon y\in{\mathds{F}}_{3}\}$ and $U=\{(1,0)\}$. Now $X$ has to be a primitive subset of $C(U)={\mathds{F}}_{3}\times\{0\}$, whence either $X=\{(-1,0)\}$ or $X=\{(1,0)\}$. But these two cases are ruled out by ()()()1.2 and ()()()1.2, respectively. Altogether, the primitive subsets of ${\mathds{F}}_{3}^{2}$ are exactly the lines avoiding the origin.
Next we look at $V={\mathds{F}}_{3}^{3}$ and $H=\{1\}\times{\mathds{F}}_{3}^{2}$, say. The case $\dim(U)=1$ is easily dismissed as in the previous paragraph, but when $U$ is a single point we can take $X=-U$ without violating ()()()1.2. This yields primitive sets such as
$P=\{(1,0,1),(1,1,-1),(1,1,0),(1,1,1),(-1,0,0)\}\,.$
As the definition of $(H,U)$-halves involves four binary choices when $\dim(H/U)=2$, one might think at first that there also are some ‘essentially different’ primitive sets of this type. It could be shown by means of a simple case analysis, however, that all five-element primitive subsets of ${\mathds{F}}_{3}^{3}$ are ‘isomorphic’ to $P$, i.e., that they are in the orbit of $P$ with respect to the automorphism group of ${\mathds{F}}_{3}^{3}$.
The configuration $P$ was first found by Lev [VL05], who observed that for $n\geq 3$ the maximal sum-free set $P\times{\mathds{F}}_{3}^{n-3}$ cannot be covered by a hyperplane, which shows that the bound $5\cdot 3^{n-3}$ in his aforementioned result is optimal. He further observed that for $n\geq 3$, $\dim(U)=0$, and ${X=-U}$ Definition 1.2()1.2 yields an aperiodic maximal sum-free subset of ${\mathds{F}}_{3}^{n}$ whose size is $\frac{1}{2}(3^{n-1}+1)$.
In general, it turns out that the cardinality of a primitive set only depends on the dimension of its period space. Before making this precise we would like to recall that for a nonempty subset $A$ of an abelian group $G$ its symmetry group is defined by
$\mathrm{Sym}(A)=\{x\in G\colon A+x=A\}\,.$
If $G$ is a ternary vector space, then $\mathrm{Sym}(A)$ is a linear subspace of $G$, and $A$ is aperiodic if and only if $\mathrm{Sym}(A)=\{0\}$. As we shall see in the next section, a simple induction along the recursive formation of primitive sets reveals the following.
Lemma 1.3.
Every primitive subset $A$ of a finite ternary vector space $V$ is maximal sum-free and satisfies $|A|=\tfrac{1}{6}(|V|+3|\mathrm{Sym}(A)|)$.
This provides enough context for the main result of this article, which we would like to state next.
Theorem 1.4.
A subset $A$ of a finite ternary vector space $V$ with $|A|>\frac{1}{6}|V|$ is maximal sum-free if and only if it is primitive.
Corollary 1.5 (Lev’s periodicity conjecture).
Let $A\subseteq{\mathds{F}}_{3}^{n}$ be a maximal sum-free set. If $A$ is aperiodic, then $n\neq 2$ and $|A|\leq\frac{1}{2}(3^{n-1}+1)$. ∎
Let us finally point out that such problems can also be studied in ${\mathds{F}}_{p}^{n}$ for $p>3$ (see, e.g., [VL23, LV23, RZ24a, RZ24b, RS71] and the concluding remarks).
2. Properties of primitive sets
In this section we collect some simple properties of primitive sets $A$, many of which are shown by unravelling their recursive definition. In such situations it will be convenient to say that $A$ is derived from $(H,U,W,X)$ if $H$, $U$, $W$, $X$ are as described in Definition 1.2()1.2. Working towards Lemma 1.3 we first look at symmetry sets.
Lemma 2.1.
If a primitive set $A$ is derived from $(H,U,W,X)$, then
$\mathrm{Sym}(A)=\mathrm{Sym}(X)\subseteq[U]\,.$
Proof.
Suppose first that $v\in\mathrm{Sym}(A)$. Since $A\cap[H]=\varnothing$, we have $v\in[H]$. For this reason, $v$ can be expressed as a difference of two vectors from $A\cap(-H)=X\cap(-U)$ and, in particular, we have $v\in[U]$. This entails $W+v=W$ and $v\in\mathrm{Sym}(X)$ follows.
Now suppose conversely that some vector $v\in\mathrm{Sym}(X)$ is given. Owing to $X\subseteq U\cup(-U)$ we have $v\in[U]$, whence $W=W+v$ and $v\in\mathrm{Sym}(A)$. ∎
We proceed with an observation of Lev (cf. [VL05]*Lemma 5).
Lemma 2.2 (Lev).
Every subset $A$ of a ternary vector space $V$ with $|A|>\frac{1}{3}|V|$ satisfies $A-A=V$.
Proof.
Let $v\in V$ be arbitrary. For cardinality reasons, the sets $A$, $A+v$, and $A-v$ cannot be disjoint. Taking into account that $-v=2v$ this yields $v\in A-A$. ∎
Proof of Lemma 1.3.
When $A$ is a hyperplane we have $|V|=3|A|$ as well as $\mathrm{Sym}(A)=[A]$, and everything is clear. We shall now establish the general case by induction on $n=\dim(V)$. In the induction step we may assume that $A$ is derived from $(H,U,W,X)$. The induction hypothesis and Lemma 2.1 yield
$|A|=|W|+|X|=\tfrac{1}{2}(|H|-|U|)+\tfrac{1}{6}(|C(U)|+3|\mathrm{Sym}(A)|)=\tfrac{1}{6}(|V|+3|\mathrm{Sym}(A)|)\,.$
Furthermore, $A$ is easily seen to be sum-free and it remains to show that there is no vector $v\in V\smallsetminus A$ such that $A\mathbin{\mathchoice{\leavevmode\vtop{ \halign{\hfil$\m@th\displaystyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\textstyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\scriptstyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\scriptscriptstyle#$\hfil\cr\cup\cr\cdot\crcr}}}}\{v\}$ is still sum-free. Assume for the sake of contradiction that some such vector $v$ existed. The fact that $X$ is, by induction, maximal sum-free in $C(U)$ implies $v\not\in C(U)$.
So $v$ belongs to one of the five sets $W$, $-W$, $(-W)+(-U)$, $W+U$, or $[H]$. The first two cases are impossible due to $W\subseteq A$ and $-W\subseteq W+W\subseteq A+A$. Next we observe that Definition 1.2()1.2()()()1.2 yields some vector $y\in(-U)\cap X$. Now $(-W)+(-U)\subseteq y-A\subseteq A-A$ and $W+U\subseteq A-y\subseteq A-A$ rule out the third and fourth possibility. Altogether this proves $v\in[H]$, which in view of Lemma 2.1 and $v\not\in A-A$ yields $|A\cap H|\leq\frac{1}{3}|H|$.
But $|A\cap H|=|W|+|U\cap X|$ and unless $\dim(H/U)=1$ we already have
$|W|=\tfrac{1}{2}(|H|-|U|)>\tfrac{1}{3}|H|\,.$
This tells us $\dim(H/U)=1$ and $U\cap X=\varnothing$. Now the second and third clause of Definition 1.2()1.2 lead to $X\subsetneqq(-U)$, which contradicts the fact that $X$ is a maximal sum-free subset of $C(U)$. ∎
Here is a rather surprising property of primitive sets, foreshadowed by [VL05]*Lemma 4, which will later be transferred to dense sum-free sets (see Proposition 4.5).
Lemma 2.3.
If $A$ is a primitive subset of a ternary vector space $V$, then $4A=A+A+A+A$ does not contain the origin.
Proof.
If $A$ is a hyperplane with $0\not\in A$, then $4A=A$ and the result is clear. Arguing again by induction on $\dim(V)$ we suppose that $A$ is derived from $(H,U,W,X)$ and consider the canonical projection $\pi\colon V\longrightarrow V/[H]\cong{\mathds{F}}_{3}$ sending $H$ to $1$. Assume for the sake of contradiction that there are four vectors $x_{1},x_{2},x_{3},x_{4}\in A$ such that $x_{1}+x_{2}+x_{3}+x_{4}=0$.
The numbers $\pi(x_{i})$ add up to zero as well, but due to $A\cap[H]=\varnothing$ none of them is itself zero. So without loss of generality we can suppose $x_{1},x_{2}\in A\cap H$ and $x_{3},x_{4}\in A\cap(-H)=X\cap(-U)$. In the quotient space $V/U$ we thus have $x_{1}+x_{2}+U=U$ and, as $W$ is a $(H,U)$-half, this is only possible if $x_{1},x_{2}\in X\cap U$. Altogether this shows $0\in 4X$, which contradicts the induction hypothesis. ∎
Lemma 2.4.
If $A$ denotes a primitive subset of a ternary vector space $V$ which is not a hyperplane, then every hyperplane $J\subseteq V$ satisfies $|A|+|A\cap J|\leq|J|$.
Proof.
Let $A$ be derived from $(H,U,W,X)$. There are three special cases we would like to address before completing the argument by induction on $\dim(V)$.
  1. (0 )
    If $\dim(H/U)\geq 2$ and $J\neq H$, then the claim follows from
    $|A|=|W|+|X|\leq\tfrac{1}{2}(|H|-|U|)+|U|\leq\tfrac{5}{9}|H|$
    and $|A\cap J|\leq|H\cap J|+|X|\leq\frac{4}{9}|H|$.
  2. (0 )
    Suppose next that $X$ is a hyperplane in $C(U)$. By Definition 1.2()1.2 we have $X=-U$ and $\dim(H/U)\geq 2$. So we can assume $J=H$, as otherwise the previous case applies. Now we have indeed
    $|A|+|A\cap H|=2|W|+|U|=|H|\,.$
  3. (0 )
    Suppose finally that $C(U)\subseteq J$. This yields $J\neq H$ and in view of the first case we can assume $\dim(H/U)=1$. Now $C(U)$ is a hyperplane and we actually have $J=C(U)$ as well as $|A|+|A\cap J|=|U|+2|X|\leq 3|U|=|J|$.
We are now ready for the induction step, where we assume that none of the above cases applies. As $C(U)\cap J$ is either a hyperplane in $C(U)$ or empty and $X$ fails to be a hyperplane, the induction hypothesis tells us $|X|+|X\cap J|\leq|U|$, whence
$|A|+|A\cap J|\leq 2|W|+|U|=|H|\,.\qed$
Fact 2.5.
If $U\subseteq H$ are two affine ternary spaces with $\dim(H/U)\geq 2$, then every $(H,U)$-half contains an affine space of dimension $\dim(U)+1$.
Proof.
The claim is easily reduced to its special case $H={\mathds{F}}_{3}^{2}$ and $U=\{(0,0)\}$, which is easily handled by means of a quick case analysis. ∎
Lemma 2.6.
Every primitive set $A$ that fails to be a hyperplane contains an affine space whose dimension exceeds that of $\mathrm{Sym}(A)$.
Proof.
Let $A$ be derived from $(H,U,W,X)$. As long as $X$ fails to be a hyperplane Lemma 2.1 allows us to appeal to an induction on the dimension of the ambient ternary space. On the other hand, if $X$ is a hyperplane, we can conclude $X=-U$ and $\dim(H/U)\geq 2$ from Definition 1.2()1.2. But now Fact 2.5 yields the desired affine space within $W$. ∎
The proof of Theorem 1.4 will proceed by induction on the dimension of the ambient space $V$. Since the intersection of a maximal sum-free subset of $V$ with a proper subspace $J$ of $V$ is in general only a sum-free subset of $J$ but not necessarily a maximal one (even if this intersection is very dense), we need to study subsets of primitive sets as well. For brevity we shall call them subprimitive in the sequel. Utilising the previous lemma one can show that sufficiently large subprimitive sets have density at least $\frac{5}{6}$ in certain affine spaces.
Lemma 2.7.
Let $B$ be a subprimitive subset of a ternary vector space $V$ which is not contained in a hyperplane. If $|B|>\frac{1}{6}(|V|+3^{k-1})$ holds for some positive integer $k$, then there exists a $k$-dimensional affine subspace $E$ of $V$ such that $|B\cap E|\geq\frac{1}{6}(5|E|+3)$.
Proof.
Let $A$ be a primitive superset of $B$. In view of
$\tfrac{1}{6}(|V|+3|\mathrm{Sym}(A)|)=|A|\geq|B|>\tfrac{1}{6}(|V|+3^{k-1})$
the dimension $\varrho$ of $\mathrm{Sym}(A)$ is at least $k-1$. Since $A$ cannot be a hyperplane, Lemma 2.6 yields a $(\varrho+1)$-dimensional affine space $Q\subseteq A$. Because of
$|A\smallsetminus B|<\tfrac{1}{6}(|V|+|Q|)-\tfrac{1}{6}|V|=\tfrac{1}{6}|Q|$
we have
$|B\cap Q|\geq|Q|-|A\smallsetminus B|>\tfrac{5}{6}|Q|\,.$
As $Q$ can be expressed as a disjoint union of $k$-dimensional affine spaces, this leads to some such space $E$ satisfying $|B\cap E|>\frac{5}{6}|E|$. Due to $|E|\equiv 3\pmod{6}$ we actually have $|B\cap E|\geq\frac{1}{6}(5|E|+3)$. ∎
Lemma 2.8.
Let $A$ be a primitive subset of a ternary vector space $V$, but not a hyperplane. If the subset $B\subseteq A$ satisfies $|B|>\frac{1}{6}|V|$ and the hyperplane $J$ is disjoint to $B$, then $J\cap A=\varnothing$ holds as well.
Proof.
Setting $T=\mathrm{Sym}(A)$ we have
$|A\smallsetminus B|<\tfrac{1}{6}(|V|+3|T|)-\tfrac{1}{6}|V|=\tfrac{1}{2}|T|\,,$
for which reason every translate $T^{\prime}$ of $T$ contained in $A$ satisfies $|T\cap B|>\frac{1}{2}|T|$. This takes immediate care of the case $[T]\subseteq[J]$, where $J$ is a union of translates of $T$. On the other hand, if $[T]\not\subseteq[J]$, then $|T^{\prime}\cap J|=\frac{1}{3}|T|$ holds for every translate $T^{\prime}$ of $T$. So, if $A$ consists of at least two such translates, we reach the contradiction $|A\smallsetminus B|\geq|A\cap J|\geq\frac{2}{3}|T|$. But, as $A$ fails to be a hyperplane, it cannot be a single translate of $T$ either. ∎
3. Inductive arguments
Let us recapitulate a fundamental result of additive combinatorics due to Kneser [Kn53, Kn55].
Theorem 3.1 (Kneser).
If $A$, $B$ are two finite nonempty subsets of an abelian group $G$ and $K=\mathrm{Sym}(A+B)$, then
$|A+B|\geq|A+K|+|B+K|-|K|\,.$
If $G$ is finite, then $|G|$ and the right side are divisible by $|K|$; this immediately implies the following well-known fact.
Corollary 3.2.
If two subsets $A$, $B$ of a finite abelian group $G$ satisfy $|A|+|B|>|G|$, then $A+B=G$.
Our other applications of Kneser’s theorem factorise through the the following statement.
Lemma 3.3.
Let $A$, $B$, $C$ be three subsets of a finite abelian group $G$ such that $A+B$ is disjoint to $C$. If $2|A|+2|B|+|C|>2|G|$ and $C\neq\varnothing$, then there is a subgroup $K$ of $G$ such that $|A+K|+|B+K|=|G|$ and $C$ is contained in a single coset of $K$.
Proof.
Notice that $A=B=\varnothing$ would yield the contradiction $|C|>2|G|$. If exactly one of $A$ and $B$ is empty, we can simply take $K=G$. So we may assume from now on that $A$ and $B$ are nonempty. It will eventually turn out that $K=\mathrm{Sym}(A+B)$ is as required. Since $A+B$ is a union of cosets of $K$, it is disjoint to $C+K$ and, therefore, Kneser’s theorem yields
$|G|\geq|A+B|+|C+K|\geq|A+K|+|B+K|+|C+K|-|K|\,,$ (3.1)
whence
$|G|\geq|A+K|+|B+K|\,.$ (3.2)
Furthermore, (3.1) entails
$2|A+K|+2|B+K|+|C+K|$ $\geq 2|A|+2|B|+|C|>2|G|$
$\geq|G|+|A+K|+|B+K|+|C+K|-|K|\,,$
i.e., $|A+K|+|B+K|>|G|-|K|$. As both sides are multiples of $|K|$, this shows in combination with (3.2) that $|A+K|+|B+K|=|G|$. In view of (3.1) we can conclude $|C+K|\leq|K|$ and, consequently, $C$ is indeed contained in a single coset of $K$. ∎
The foregoing lemma will be used to handle a major case in our inductive proof of Theorem 1.4. We will denote the statement that a given ternary vector space satisfies the induction hypothesis by $\Phi(V)$. More precisely, $\Phi(V)$ means that if $V^{\prime}$ refers to a ternary vector space with $\dim(V^{\prime})<\dim(V)$, then every sum-free set $A^{\prime}\subseteq V^{\prime}$ with $|A^{\prime}|>\frac{1}{6}|V^{\prime}|$ is subprimitive. Since Theorem 1.4 is clear when $\dim(V)\leq 2$, we will know that $\Phi({\mathds{F}}_{3}^{3})$ is true when looking at the three-dimensional case later.
Proposition 3.4.
Given a finite ternary vector space $V$ satisfying $\Phi(V)$, let $A$ be a sum-free subset and $H$ a hyperplane not containing the origin. If $|A|>\frac{1}{6}|V|$, $[H]\cap A=\varnothing$, and $\mathrm{aff}(A\cap(-H))$ is a proper affine subspace of $-H$, then $A$ is subprimitive.
Proof.
Due to the last assumption $U=\mathrm{aff}(H\cap(-A))$ is a proper affine subspace of $H$. If $U$ should be empty, we have $A\subseteq H$ and the claim is clear. So we can henceforth suppose ${U\neq\varnothing}$, whence $A\cap(-U)\neq\varnothing$. We shall show that there exist an $(H,U)$-half $W$ and a primitive subset $X$ of $C(U)$ such that $A$ is contained in $W\mathbin{\mathchoice{\leavevmode\vtop{ \halign{\hfil$\m@th\displaystyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\textstyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\scriptstyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\scriptscriptstyle#$\hfil\cr\cup\cr\cdot\crcr}}}}X$ —a set which is either primitive or contained in a hyperplane.
To this end we set $A_{y}=A\cap(U+y)$ for every $y\in[H]\smallsetminus[U]$. Since $A_{y}+A_{-y}\subseteq-U$ cannot hold with equality, Corollary 3.2 informs us that
$|A_{y}|+|A_{-y}|\leq|U|\,.$ (3.3)
By summing this over an appropriate half-system of representatives of $[H]/[U]$ we infer $|A\cap(H\smallsetminus U)|\leq\frac{1}{2}(|H|-|U|)$, which in turn leads to
$|A\cap C(U)|>\tfrac{1}{6}|V|-\tfrac{1}{2}(|H|-|U|)=\tfrac{1}{6}|C(U)|\,.$
Thus the inductive hypothesis $\Phi(V)$ yields a primitive subset $X\subseteq C(U)$ covering $A\cap C(U)$. If $X$ is a hyperplane, then it needs to be equal to $-U$ and
$X\cap[H]=\varnothing$ (3.4)
follows. On the other hand, if $X$ is not a hyperplane, then Lemma 2.8 applied to $A\cap C(U)$ shows that (3.4) is still true.
Claim 3.5.
For every $z\in[H]\smallsetminus[U]$ one of the sets $A_{z}$, $A_{-z}$ is empty.
Proof.
It follows from (3.3) that
$|A\cap(H\smallsetminus U)|\leq\tfrac{1}{2}(|H|-3|U|)+|A_{z}|+|A_{-z}|\,.$
In view of $A\subseteq(H\smallsetminus U)\mathbin{\mathchoice{\leavevmode\vtop{ \halign{\hfil$\m@th\displaystyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\textstyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\scriptstyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\scriptscriptstyle#$\hfil\cr\cup\cr\cdot\crcr}}}}C(U)$ and $|A|>\frac{1}{6}|V|$ this yields
$\tfrac{3}{2}|U|<|A_{z}|+|A_{-z}|+|A\cap C(U)|\,,$ (3.5)
Next we contend that
$|A\cap C(U)|+|A\cap U|\leq|U|\,.$ (3.6)
If $A\cap C(U)$ fails to be a hyperplane, this can be seen by applying Lemma 2.4 to $C(U)$, $X$ here in place of $V$, $A$ there. In the remaining case we have $X=-U$ and (3.6) is obvious.
By doubling (3.5) and taking (3.6) into account we learn
$2|U|<2|A_{z}|+2|A_{-z}|+|A\cap(-U)|\,.$
Now Lemma 3.3 shows that there is a subspace $K$ of $[U]$ such that $|A_{z}+K|+|A_{-z}+K|=|U|$ and $A\cap(-U)$ is contained in a translate of $K$. Appealing to $\mathrm{aff}((-U)\cap A)=-U$ we infer $K=[U]$ and, therefore, one of $A_{z}$, $A_{-z}$ has to be empty. ∎
As the claim demonstrates, there is an $(H,U)$-half $W$ covering $A\cap(H\smallsetminus U)$. Clearly, if Definition 1.2()1.2()()()1.2 is satisfied, then $A$ is contained in the primitive set derived from $(H,U,W,X)$. It remains to treat the exceptional case that $\dim(H/U)=1$ and $X=-U$. Now
$W\mathbin{\mathchoice{\leavevmode\vtop{ \halign{\hfil$\m@th\displaystyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\textstyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\scriptstyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\scriptscriptstyle#$\hfil\cr\cup\cr\cdot\crcr}}}}(-U)\mathbin{\mathchoice{\leavevmode\vtop{ \halign{\hfil$\m@th\displaystyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\textstyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\scriptstyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\scriptscriptstyle#$\hfil\cr\cup\cr\cdot\crcr}}}}(U+(-W))$
is a hyperplane in $V$ covering $A$, so that $A$ is again subprimitive. ∎
The three-dimensional case of Theorem 1.4 was first proved by Lev (see [VL05]*Lemma 3) using an exhaustive case analysis. For the sake of completeness we would like to mention an alternative argument based on Proposition 3.4.
Corollary 3.6 (Lev).
Every sum-free set $A\subseteq{\mathds{F}}_{3}$ with $|A|\geq 5$ is subprimitive and contains a line.
Proof.
We shall show first that every five-element sum-free set $A\subseteq{\mathds{F}}_{3}$ contains four coplanar points. To this end we observe that thirteen lines of ${\mathds{F}}_{3}^{3}$ pass through the origin. Five of them intersect $A$, while the eight other ones do not. There are $\binom{5}{2}>8$ pairs of points from $A$ and each pair determines the direction of a line. Thus there are either two parallel lines intersecting $A$ at least twice, or there is a line intersecting $A$ thrice. In both cases, there are indeed four coplanar points in $A$.
Now let $A\subseteq{\mathds{F}}_{3}$ be an arbitrary sum-free set with $|A|\geq 5$. If the entire set is coplanar, then it is clearly subprimitive and a quick case analysis discloses that $A$ contains three collinear points. Otherwise, there is a subset $A^{\prime}\subseteq A$ of size five which fails to be coplanar. By the first paragraph there is a plane $H$ with $|A^{\prime}\cap H|=4$, Proposition 3.4 shows that $A^{\prime}$ is primitive, and Lemma 1.3 yields $A^{\prime}=A$. Finally, by Lemma 2.6 $A$ contains a line. ∎
Here is a straightforward consequence, whose easy proof we omit.
Corollary 3.7.
If $A\subseteq{\mathds{F}}_{3}^{3}$ is a subprimitive set of size four, then either $A$ is contained in a plane or one point in $A$ is the sum of the three other ones. ∎
A simple averaging arguments transfers the last part of Corollary 3.6 to higher dimensions.
Corollary 3.8.
For every at least three-dimensional ternary vector space $V$, every sum-free set $A\subseteq V$ with $|A|>\frac{1}{6}|V|$ contains a line.
Proof.
By averaging there exists a three-dimensional subspace $L\leq V$ such that
$|A\cap L|\geq\frac{26|A|}{|V|-1}>\frac{26}{6}>4\,,$
i.e., $|A\cap L|\geq 5$. So by Corollary 3.6 $A\cap L$ contains a line. ∎
On first reading the next result might look like a variant of Proposition 3.4. However, as it is quite unrelated to the recursive structure of primitive sets, it lies much closer to the surface.
Proposition 3.9.
Let $A$ be a sum-free subset of a finite ternary vector space $V$ of size $|A|>\frac{1}{6}|V|$. If there exists a hyperplane $H$ not passing through the origin such that $A\cap H=\varnothing$ and $\mathrm{aff}(A\cap[H])\neq[H]$, then $A$ is subprimitive.
Proof.
Let $U$ be a hyperplane in $[H]$ covering $\mathrm{aff}(A\cap[H])$. If $|A\cap(-H)|>\frac{1}{3}|H|$, then Lemma 2.2 yields $A\cap[H]=\varnothing$, whence $A$ is covered by the hyperplane $-H$. Suppose now that $|A\cap(-H)|\leq\frac{1}{3}|H|$, wherefore
$|A\cap U|>\tfrac{1}{6}|H|=\tfrac{1}{2}|U|\,.$ (3.7)
In particular, $U$ does not contain the origin and there exists a partition of $-H$ into three translates $X$, $Y$, $Z$ of $U$ such that $Y=X+U$ and $Z=Y+U$. For reasons of cyclic symmetry we can suppose
$|A\cap Z|=\min\{|A\cap X|,|A\cap Y|,|A\cap Z|\}\,.$ (3.8)
If $|A\cap X|+|A\cap U|>|U|$, then $A$ is disjoint to $X\pm U$, whence $A$ is covered by the hyperplane containing $X$ and $U$. So we may assume from now on that
[…]
and, similarly, that $|A\cap X|>\frac{1}{4}|U|$. In particular, neither $A\cap X$ nor $A\cap Y$ is empty. Together with
$2|A\cap X|+2|A\cap U|+|A\cap Y|$ $=|A|+|A\cap U|+(|A\cap X|-|A\cap Z|)$
[…]
and Lemma 3.3 this tells us that there is a subspace $K$ of $[U]$ such that
$|(A\cap X)+K|+|(A\cap U)+K|=|U|$
and $A\cap Y$ is covered by a single translate of $K$. We cannot have $K=[U]$, for then one of $A\cap X$ or $A\cap U$ had to be empty. So (3.9) reveals $\dim(U/K)=1$ and we see that $A\cap X$ is contained in a single translate of $K$, while $A\cap U$ is covered by two translates of $K$.
Let $\psi\colon V\longrightarrow{\mathds{F}}_{3}^{3}$ be the epimorphism with kernel $K$ satisfying
$\psi[A\cap U]=\{(0,1,1),(0,1,2)\}\quad\text{ and }\quad\psi[A\cap X]=\{(1,0,0)\}\,.$
Since $|A\cap Z|\leq|A\cap Y|\leq|K|$, we have
$|A\cap X|+|A\cap U|>|A|-2|K|>2|K|$
and, hence,
$|A\cap\psi^{-1}(1,0,0)|+|A\cap\psi^{-1}(0,1,1)|>|K|\,.$
By Corollary 3.2 this yields $(1,0,0)\pm(0,1,1)\not\in\psi[A]$ and, similarly, $(1,0,0)\pm(0,1,2)\not\in\psi[A]$ holds as well. This leaves us with $\psi[A\cap Y]=\{(1,1,0)\}$ and $\psi[A\cap Z]=\{(1,2,0)\}$, so that altogether
$\psi[A]=\{(0,1,1),(0,1,2),(1,0,0),(1,1,0),(1,2,0)\}\,.$
As this subset of ${\mathds{F}}_{3}^{3}$ is primitive, $A$ is indeed subprimitive. ∎
Both propositions of this section combine as follows.
Conclusion 3.10.
Assume $\Phi({\mathds{F}}_{3}^{n})$, where $n\geq 2$, and set $V_{ij}=\{(i,j)\}\times{\mathds{F}}_{3}^{n-2}$ for every point $(i,j)\in{\mathds{F}}_{3}^{2}$. Let $A\subseteq{\mathds{F}}_{3}^{n}$ be a sum-free set with ${|A|>\tfrac{1}{2}\cdot 3^{n-1}}$ and $|A\cap V_{01}|>\tfrac{1}{2}\cdot 3^{n-2}$. If for every $i\in{\mathds{F}}_{3}$ at least one of the sets $A\cap V_{1i}$, $A\cap V_{2,1-i}$ is empty, then $A$ is subprimitive.
Proof.
Set $A_{ij}=A\cap V_{ij}$ for every $(i,j)\in{\mathds{F}}_{3}^{2}$. Corollary 3.2 yields $A_{01}-A_{01}=V_{00}$ and $A_{01}+A_{01}=V_{02}$, whence $A_{00}=A_{02}=\varnothing$. By symmetry we can assume further that
  1. (0 )
    either $A_{10}=A_{11}=A_{12}=\varnothing$
  2. (0 )
    or $A_{10}=A_{12}=A_{20}=\varnothing$.
If (0 ) holds, then the hyperplane $H=V_{10}\mathbin{\mathchoice{\leavevmode\vtop{ \halign{\hfil$\m@th\displaystyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\textstyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\scriptstyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\scriptscriptstyle#$\hfil\cr\cup\cr\cdot\crcr}}}}V_{11}\mathbin{\mathchoice{\leavevmode\vtop{ \halign{\hfil$\m@th\displaystyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\textstyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\scriptstyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\scriptscriptstyle#$\hfil\cr\cup\cr\cdot\crcr}}}}V_{12}$ fulfils the hypothesis of Proposition 3.9 and $A$ is indeed subprimitive. Similarly, in case (0 ) we can apply Proposition 3.4 to the hyperplane $H=V_{01}\mathbin{\mathchoice{\leavevmode\vtop{ \halign{\hfil$\m@th\displaystyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\textstyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\scriptstyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\scriptscriptstyle#$\hfil\cr\cup\cr\cdot\crcr}}}}V_{11}\mathbin{\mathchoice{\leavevmode\vtop{ \halign{\hfil$\m@th\displaystyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\textstyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\scriptstyle#$\hfil\cr\cup\cr\cdot\crcr}}}{\leavevmode\vtop{ \halign{\hfil$\m@th\scriptscriptstyle#$\hfil\cr\cup\cr\cdot\crcr}}}}V_{21}$. ∎
4. Quadruple sums
Summarising our current state of knowledge, we have shown the ‘easy’ direction of Theorem 1.4 for all ternary spaces $V$, and we know that the ‘difficult’ direction holds in the low-dimensional cases $\dim(V)\leq 3$ (cf. Corollary 3.6), which gives the induction hypothesis $\Phi({\mathds{F}}_{3}^{4})$. Before we can complete the general induction step, however, we need to study the case $n=4$ separately (cf. Lemma 4.3). Roughly speaking, this is because that case is required for establishing that dense sum-free sets $A$ satisfy $0\not\in 4A$.
We would like to point out that, as reported in [VL05], Lev verified the case $n=4$ of Corollary 1.5 electronically. It seems reasonable to believe that the programme he wrote could confirm Lemma 4.3 as well. Nevertheless, we did not pursue this direction here. Rather, we argue that the results from the previous section allow us to split this problem into a manageable number of cases. We commence with sets containing two parallel lines.
Lemma 4.1.
If a sum-free set $A\subseteq{\mathds{F}}_{3}^{4}$ with $|A|\geq 14$ contains two parallel lines, then it is subprimitive.
Proof.
Arguing indirectly we assume that $A$ is not subprimitive. For every $k\in\{1,2,3,4\}$ let $\pi_{k}\colon{\mathds{F}}_{3}^{4}\longrightarrow{\mathds{F}}_{3}$ be the projection onto the $k^{\mathrm{th}}$ coordinate. Set $V_{ij}=\{(i,j)\}\times{\mathds{F}}_{3}^{2}$ for every point $(i,j)\in{\mathds{F}}_{3}^{2}$. Pick a two-dimensional affine space $Q\subseteq{\mathds{F}}_{3}^{4}$ such that $A\cap Q$ contains two parallel lines and, subject to this, $|A\cap Q|$ is maximal. Due to $|A\cap Q|\geq 6>3$ the origin cannot be in $Q$. So an appropriate choice of coordinates allows us to assume $Q=V_{01}$ and that $B_{3}=\{(0,1,y,z)\in{\mathds{F}}_{3}^{4}\colon y\neq 0\}$ is a subset of $A$.
Claim 4.2.
If two points $p,q\in A$ satisfy $\pi_{1}(p)=\pi_{1}(q)\neq 0$, then $\pi_{3}(p)=\pi_{3}(q)$.
Proof.
Assume for the sake of contradiction that this fails, i.e., that there are $p,q\in A$ such that $\pi_{1}(p)=\pi_{1}(q)\neq 0$, but $\pi_{3}(p)\neq\pi_{3}(q)$. Owing to ±$\pm(p-q)\not\in B_{3}$ we have $\pi_{2}(p)=\pi_{2}(q)$. Without loss generality we can assume that $p,q\in V_{11}$. We intend to show that $H=\pi_{2}^{-1}(1)$ satisfies the assumptions of Proposition 3.4. Due to $B_{3}-B_{3}=V_{00}$, $\{p,q\}-B_{3}=V_{10}$, and $B_{3}-\{p,q\}=V_{20}$ we have indeed $A\cap[H]=\varnothing$. Moreover $B_{3}+B_{3}=V_{02}$ and $B_{3}+\{p,q\}=V_{12}$ show that $A\cap(-H)$ is contained in the affine space $V_{22}$. ∎
By Conclusion 3.10 there exists some $i\in{\mathds{F}}_{3}$ for which neither $A\cap V_{1i}$ nor $A\cap V_{2,1-i}$ is empty. Pick $x\in A\cap V_{1i}$, $y\in A\cap V_{2,1-i}$ and set $r_{3}=\pi_{3}(x)$, $s_{3}=\pi_{3}(y)$. As $x+y\not\in V_{00}\smallsetminus B_{3}$, we have $r_{3}+s_{3}=0$. By Claim 4.2 all points $x^{\prime}\in A$ with $\pi_{1}(x^{\prime})=1$ satisfy $\pi_{3}(x^{\prime})=r_{3}$ and, similarly, all $y^{\prime}\in A$ with $\pi_{1}(y^{\prime})=2$ have the property $\pi_{3}(y^{\prime})=s_{3}=2r_{4}$. Consequently, $A\smallsetminus B_{3}$ is contained in the three-dimensional vector space $W_{3}=\{z\in{\mathds{F}}_{3}^{4}\colon\pi_{3}(z)=r_{3}\pi_{1}(z)\}$.
In view of $|A\smallsetminus B_{3}|=8>5$ it follows that some two-dimensional affine space covers $A\smallsetminus B_{3}$. By our extremal choice of $Q=V_{01}$ this yields $|A\cap V_{01}|\geq 8$ and we can assume that the points $(0,1,0,\pm 1)$ are in $A$, which causes $B_{4}=\{(0,1,y,z)\in{\mathds{F}}_{3}^{4}\colon z\neq 0\}$ to be a subset of $A$. By exchanging the rôles of the third and fourth coordinate in the arguments presented to far we see that $A\smallsetminus B_{4}$ lives in a three-dimensional space of the form $W_{4}=\{z\in{\mathds{F}}_{3}^{4}\colon\pi_{4}(z)=r_{4}\pi_{1}(z)\}$. But now $\dim(W_{3}\cap W_{4})=2$ yields $|A\cap W_{3}\cap W_{4}|\leq 3$ and we are led to the contradiction $|A|=|B_{3}\cup B_{4}|+|A\cap W_{3}\cap W_{4}|\leq 11$. ∎
Lemma 4.3.
Every sum-free set $A\subseteq{\mathds{F}}_{3}^{4}$ with $|A|\geq 14$ is subprimitive.
Proof.
Let $\psi\colon{\mathds{F}}_{3}^{4}\longrightarrow{\mathds{F}}_{3}^{3}$ be the projection deleting the last coordinate. Since $A$ contains a line (by Corollary 3.8), we may assume that there exists a point $p\in{\mathds{F}}_{3}^{3}$ such that $\psi^{-1}(p)\subseteq A$. Set $B=\{q\in{\mathds{F}}_{3}^{3}\colon|\psi^{-1}(q)\cap A|\geq 2\}$ and $C=\{q\in{\mathds{F}}_{3}^{3}\colon\psi^{-1}(q)\cap A\neq\varnothing\}$. As $A$ is sum-free, we have
$(B+B)\cap C=(B-B)\cap C=\varnothing\,.$ (4.1)
In particular, $B$ is sum-free. If there exists a point $q\neq p$ in ${\mathds{F}}_{3}^{3}$ with $\psi^{-1}(q)\subseteq A$, then $A$ contains two parallel lines and Lemma 4.1 tells us that $A$ is indeed subprimitive. So we may assume from now on that $p$ is the only such point. In view of $|A|\geq 14$ this implies
$|B|+|C|\geq 13\,.$ (4.2)
Moreover, $C$, $C+p$, and $C-p$ are disjoint subsets of ${\mathds{F}}_{3}^{3}$, whence
$|C|\leq 9\,.$ (4.3)
Claim 4.4.
There are four coplanar points in $B$.
Proof.
If $|B|\geq 5$ this follows from the already established case $V={\mathds{F}}_{3}^{3}$ of Theorem 1.4. Due to (4.2) and (4.3) it remains to discuss the case that $|B|=4$ and $|C|=9$. Take an arbitrary point $c\in C\smallsetminus B$. Using (4.1) it is easy to show that $B\cup\{c\}$ is sum-free. Hence, $B$ is subprimitive and either we are done by Corollary 3.7 or one point of $B$ is the sum of the three other ones. For reasons of symmetry we may assume that
$B=\{(1,0,0),(0,1,0),(0,0,1),(1,1,1)\}\,.$
But now (4.1) entails
$C\subseteq B\cup\{(1,2,2),(2,1,2),(2,2,1)\}\,,$
which contradicts $|C|=9$. ∎
Let $H\subseteq{\mathds{F}}_{3}^{3}$ be a plane intersecting $B$ in at least four points. This plane cannot contain the origin, Lemma 2.2 yields $B-B=[H]$, and Kneser’s theorem implies $|(B+B)\cap(-H)|\geq 7$. Hence, (4.1) tells us $C\cap[H]=\varnothing$ and that $C\cap(-H)$ can be covered by a line. So by Proposition 3.4 applied to $\psi^{-1}[H]$ the set $A$ is indeed subprimitive. ∎
Now we keep a promise given in Section 2.
Proposition 4.5.
If a sum-free subset $A$ of a ternary space $V$ satisfies $|A|>\frac{1}{6}|V|$, then ${0\not\in 4A}$.
Proof.
Assume for the sake of contradiction that $x_{1}$, $x_{2}$, $x_{3}$, and $x_{4}$ are four elements of $A$ summing up to 00, and let $L$ be the subspace of $V$ generated by these vectors. It is easily seen that $L$ is three-dimensional. Moreover, $\{x_{1},x_{2},x_{3},x_{4}\}$ is a maximal sum-free subset of $L$, whence $|A\cap L|=4$.
Let us now consider any four-dimensional subspace $Q$ of $V$ that contains $L$. By Lemma 2.3 $A\cap Q$ cannot be subprimitive in $Q$ and, therefore, Lemma 4.3 entails $|A\cap Q|\leq 13$. This proves $|A\cap(Q\smallsetminus L)|\leq 9$ and by averaging over $Q$ we infer
$|A|\leq|A\cap L|+\tfrac{9}{54}|V\smallsetminus L|=\tfrac{1}{6}(|L|-3)+\tfrac{1}{6}(|V|-|L|)<\tfrac{1}{6}|V|\,,$
which is absurd. ∎
5. Proof of the main result
It turns out that Conclusion 3.10 allows us to prove the main theorem by induction, because there always exists a coordinate system with respect to which its hypotheses are valid. For the first of them this is shown as follows.
Proposition 5.1.
Let $V$ be an at least three-dimensional ternary vector space satisfying $\Phi(V)$. For every sum-free set $A\subseteq V$ with $|A|>\frac{1}{6}|V|$ there is an affine subspace $Q$ of $V$ such that $\dim(V/Q)=2$ and $|A\cap Q|\geq\frac{1}{2}(|Q|+3)$.
Proof.
Recall that $A$ contains a line by Corollary 3.8. In other words, there is a one-dimensional affine subspace $L$ such that $|A\cap L|=\frac{1}{6}(5|L|+3)$. Now let $k\leq\dim(V)-2$ be the largest integer such that $|A\cap U|\geq\frac{1}{6}(5|U|+3)$ holds for some $k$-dimensional affine space $U$. If $k=\dim(V)-2$ we are done, so suppose from now on that this is not the case. Note that $|A\cap U|>\frac{1}{2}|U|$ yields $0\not\in U$ and $|C(U)|=3|U|$. Setting $q=\frac{|V|}{9|U|}$ an averaging argument shows that there is a hyperplane $J$ such that $C(U)\leq J\leq V$ and
$|A\cap J|$ $\geq|A\cap C(U)|+\frac{|J|-|C(U)|}{|V|-|C(U)|}|A\smallsetminus C(U)|=\frac{q-1}{3q-1}|A|+\frac{2q}{3q-1}|A\cap C(U)|$
$>\frac{|J|}{6}+\frac{5q|U|-|J|}{9q-3}=\frac{|J|}{6}+\frac{2q|U|}{9q-3}>\frac{3|J|+4|U|}{18}\,.$
In particular, $A\cap J$ is subprimitive in $J$ by our assumption $\Phi(V)$. Furthermore, if $A\cap J$ is not contained in a hyperplane of $J$, then Lemma 2.7 yields a $(k+1)$-dimensional affine subspace $U^{\prime}$ of $J$ such that $|A\cap U^{\prime}|\geq\frac{1}{6}(5|U^{\prime}|+3)$, contrary to the maximality of $k$. Hence there is a hyperplane $Q$ in $J$ covering $A\cap J$. Due to
$|A\cap Q|=|A\cap J|>\tfrac{1}{6}(|J|+4)>\tfrac{1}{2}(|Q|+1)$
we have thereby found the desired affine subspace $Q$. ∎
Proof of Theorem 1.4.
We may assume $V={\mathds{F}}_{3}^{n}$ for some $n\geq 2$ and, for reasons of induction, that $\Phi(V)$ holds. Because of Lemma 1.3 it suffices to show that every sum-free set $A\subseteq{\mathds{F}}_{3}^{n}$ of size $|A|>\frac{1}{6}\cdot 3^{n}$ is subprimitive. Let $\varphi\colon{\mathds{F}}_{3}^{n}\longrightarrow{\mathds{F}}_{3}^{2}$ be the projection onto the first two coordinates and set $V_{ij}=\varphi^{-1}(i,j)$, $A_{ij}=A\cap V_{ij}$ for every point $(i,j)\in{\mathds{F}}_{3}^{2}$. Owing to Proposition 5.1 we may assume that $|A_{01}|\geq\frac{1}{2}(3^{n-2}+3)$, which yields $A_{01}+A_{01}=V_{02}$.
On the other hand, for every $i\in{\mathds{F}}_{3}$ Proposition 4.5 discloses $A_{01}+A_{01}+A_{1i}+A_{2,1-i}\neq V_{00}$, which together with $(0,1)+(0,1)+(1,i)+(2,1-i)=(0,0)$ informs us that at least one of the sets $A_{1i}$, $A_{2,1-i}$ is empty. In the light of Conclusion 3.10 this shows that $A$ is indeed subprimitive. ∎
6. Concluding remarks
Given an arbitrary finite abelian group $G$ let $t(G)$ be the largest possible size of an aperiodic maximal sum-free set $A\subseteq G$. For some groups $G$ no such sets exist and then we put $t(G)=0$. It seems plausible that this invariant of $G$ tends to be ‘small’ if $G$ has ‘many’ subgroups and thus the special case that $G$ is a finite vector space might be especially interesting.
As pointed out in the introduction, Davydov and Tombak [DT] proved
$t({\mathds{F}}_{2}^{n})=\begin{cases}1&\text{ if $n=1$}\\ 0&\text{ if $n=2,3$}\\ 2^{n-2}+1&\text{ if $n\geq 4$}\,,\end{cases}$
while Lev’s periodicity conjecture asserts
$t({\mathds{F}}_{3}^{n})=\begin{cases}\tfrac{1}{2}(3^{n-1}+1)&\text{ if $n\neq 2$}\\ 0&\text{ if $n=2$}\,.\end{cases}$
For primes $p$ satisfying $p\equiv 1\pmod{3}$ results of Rhemtulla and Street [RS70, RS71] imply
$t({\mathds{F}}_{p}^{n})=\tfrac{1}{3}(p-1)p^{n-1}\,,$
while [RZ24a] yields
$t({\mathds{F}}_{p}^{n})=\begin{cases}\frac{1}{3}(p+1)&\text{ if }n=1\\ \frac{1}{3}(p-2)p^{n-1}&\text{ if }n\geq 2\end{cases}$
whenever $p\equiv 2\pmod{3}$ and $p\geq 11$. This leaves only the case $p=5$ open. It is not difficult to check directly that $t({\mathds{F}}_{5})=2$ and $t({\mathds{F}}^{2}_{5})=5$ (see [LV23]). Moreover, the main result of [RZ24b] entails $t({\mathds{F}}_{5}^{3})=28$. For $n\geq 4$, however, nothing seems to be known about $t({\mathds{F}}_{5}^{n})$ and nobody ever conjectured anything.